# Open Problems, and Future Trends

### v1

@@ Abstract
As an emerging technology that enables seamless integration between the physical and virtual worlds, the Metaverse has great potential to be deployed in the industrial production field with the development of extended reality (XR) and next-generation communication networks. This deployment, called the Industrial Metaverse, is used for product design, production operations, industrial quality inspection, and product testing. However, there lacks of in-depth understanding of the enabling technologies associated with the Industrial Metaverse. This encompasses both the precise industrial scenarios targeted by each technology and the potential migration of technologies developed in other domains to the industrial sector. Driven by this issue, in this article, we conduct a comprehensive survey of the state-of-the-art literature on the Industrial Metaverse. Specifically, we first analyze the advantages of the Metaverse for industrial production. Then, we review a collection of key enabling technologies of the Industrial Metaverse, including blockchain (BC), digital twin (DT), 6G, XR, and artificial intelligence (AI), and analyze how these technologies can support different aspects of industrial production. Subsequently, we present numerous formidable challenges encountered within the Industrial Metaverse, including confidentiality and security concerns, resource limitations, and interoperability constraints. Furthermore, we investigate the extant solutions devised to address them. Finally, we briefly outline several open issues and future research directions of the Industrial Metaverse.
@@
Metaverse, industrial, enabling technologies, BC, DT, 6G, XR, AI.
@@ I Introduction
The Metaverse is a nascent concept that presents a model of interconnection between the virtual and real worlds[1]. This digital system integrates both worlds into a cohesive system, providing users with a high degree of freedom to create content and modify modules as they wish, while also allowing different participants to define their own expressions. Some scholars define the Metaverse as a collection of virtual time and space, which leverages technologies such as virtual reality (VR) and augmented reality (AR). People continuously generate new content through smart wearables as inputs. It is also considered to be the latest phase of visual immersion technology, which has applications in all areas of production and life as an online digital space. Although there is no precise definition of the term “Metaverse” at this stage, its underlying technical system has matured and has been successfully applied in fields such as education, healthcare, industrial manufacturing, among others. The Industrial Metaverse refers to the establishment of a shared virtual space empowered by Metaverse technologies, to support multi-user, multi-device industrial scenarios for 3D modeling and immersive interaction.
@@
Especially in the context of the COVID-19 pandemic, which has led to a widespread shift towards remote work, and the continuous advancement of digital technologies, many production activities that traditionally required physical presence can be transformed to enable employees to operate online, similar to a digital twin (DT). However, the difference is that DT is only as a digital replica of the physical elements of the real world, realizing the mapping from the physical world to the digital world. In contrast, Industrial Metaverse is a collective virtual shared space that contains all virtual worlds and derived elements, and virtual space digital individuals for online sharing, interactive perception and content generation. This feature is particularly suited to industrial scenarios, such as detecting part performance through nearly costless fault simulations, creating experimental conditions that cannot be replicated in reality, and expanding operational capabilities through the linkage between the virtual and physical worlds. This can guide physical industrial operations, reduce costs, and improve efficiency by empowering all aspects of the industrial process. Compared to typical virtual worlds, the emphasis of an industrial virtual world lies in simulating real-world industrial processes, manufacturing environments, or operational scenarios. The Industrial Metaverse places a higher demand on cost control for industrial manufacturing, requiring its design and implementation to consider cost efficiency and optimization, aiming to minimize deployment and maintenance costs. Additionally, the Industrial Metaverse involves collaboration among different enterprises in aspects such as industrial production and supply chain management, thus requiring a high level of trust in the execution of interactions. The key benefits of Metaverse are highlighted as follows:
@@
Sure, here’s the same table without the blue font color commands:
TABLE I: Key Abbreviations.
@@
XR Extended Reality
@@
IIoT Industrial Internet of Things
XRI XR-IoT
VR Virtual Reality
AR Augmented Reality
@@
SVM Support Vector Machines
LSTM Long Short-Term Memory
GANs Generative Adversarial Networks
@@
FL Federated Learning
@@
AI Artificial Intelligence
ASP AI Service Provider
MEC Mobile Edge Computing
NFT Non-Fungible Token
EIoT Energy Internet of Things
HCI Human-computer Interaction
O&M Operations & Maintenance
BIM Building Information Modeling
PLC Product Lifecycle
CMS Cyber-Manufacturing Systems
HMDs Head-mounted Displays
DeAI Decentralized AI
CeAI Centralized AI
AIGC AI Generated Content
URLLC Ultra-reliable and Low-latency Communication
Here’s the table without the blue font color commands:
TABLE II: Survey papers of Industrial Metaverse.
Ref. Year Scope Technologies Key Contents
Chang et al.[4] 2022 Education, product testing, telecommuting, production optimization, smart cities Computer vision, DT, Network, BC Three new Metaverse architectures at the edge, and technologies that help the Metaverse interact and share digital data.
Said et al.[7] 2023 HCI XR This work illustrates the complexity of Metaverse online learning from the perspective of HCI.
Dong et al.[9] 2023 Task-oriented communications Semantic communication Specific requirements for the three main types of tasks in the Industrial Metaverse.
Our paper 2023 Industry 5.0 (Industrial data collection and storage, product design, operation training, system manufacturing, industrial quality control, etc.) BC, DT, 6G, XR, AI An extensive survey of the Industrial Metaverse. Particularly, For the first time, we extensively discussed the role of several key enabling technologies at different stages of industry. We comprehensively analyze the main existing problems and solutions in the Industrial Metaverse. We first emphasize more comprehensive open challenges and research directions for the characteristics of industrial scenarios.
@@ I-A Related Works
Neal Stephenson’s novel “Snow Crash” was the first to introduce the concept of the Metaverse, a digital landscape constructed through AR technology. Since then, there has been an explosion in literature on the Metaverse, with much of it focused on its fundamental characteristics[10], empowering technologies[11], and privacy and security aspects. Some works also analyze the platform deployment and necessary components required for its application scenarios. Our research on the industrial applications of the Metaverse is comprehensive and detailed in Table II.
@@
Although some of the works do not specifically focus on industrial applications of the Metaverse, the enabling technologies involved can be directly transferred to the manufacturing industry. For instance, Dionisio et al. [10] provide a comprehensive summary of the four key characteristics of the Metaverse, namely, realism, ubiquity, interoperability, and scalability, which can be leveraged to improve manufacturing operations. Similarly, Wang et al. [12] and Yu et al. [13] analyze the potential of MEC in the Metaverse, and outline the basic components required to deploy MEC in the 6G network environment, which can be applied in various predictive maintenance scenarios in the manufacturing industry. Furthermore, Kye et al.[14] conduct a thorough analysis of the potential, as well as the constraints, of utilizing the Metaverse in educational contexts. This encompasses the domains of curriculum design and platform development, which could be effectively applied to skill training scenarios within the industry.
@@ I-C Structure of The Survey
The sections of this paper are structured as follows. Section I presents an overview of the fundamental principles underlying the Metaverse and elucidates the rationale behind its integration into industrial contexts. Additionally, it provides an inventory of pertinent Metaverse literature reviews, comparing them to identify the innovative aspects of our current undertaking. Section II discusses around the five core enabling technologies of the Metaverse and summarizes the industrial applications and existing problems of each technology. Section III succinctly summarizes the four primary challenges encountered in the Industrial Metaverse. In addition, in Section IV we collate some of the existing standardization efforts and platforms for the Industrial Metaverse. Moving forward, Section V offers a comprehensive compilation of various unresolved hurdles prevailing in the contemporary Industrial Metaverse. Finally, our discourse culminates with a conclusive remark in Section VI. Table I indicates the list of abbreviations employed in this paper. Fig. 2 outlines the organization of the survey.
Figure 3: Industrial Metaverse Architecture. It consists of three layers: the data input layer, the enabling layer, and the industrial application layer. The enabling layer includes six components: AI, DT, BC, XR, the Metaverse management center, and the data processing system. In this context, the data input layer is composed of various industrial IoT devices. The collected multimodal industrial data is transmitted to the data processing system for preprocessing. The processed data can be stored in the BC or streamed to the AI component for analysis. The results of data analysis are used to respond to requests from the Metaverse management center, and decision outcomes are fed back to the BC. The BC performs updates on the network and DT. The DT achieves proportional replication and synchronization of virtual and physical devices and transfers parameter information to XR for 3D modeling. XR displays decision solutions through human-machine interaction devices and is applied in various industrial scenarios.
@@ II The enabling technologies for Metaverse
This section introduces several core enabling technologies for the Industrial Metaverse, including BC, DT, XR, AI, and 6G. The primary roles of these enabling technologies in the Industrial Metaverse are as follows:
@@
If DT, XR, and AI are prerequisites for enterprise decision-making and intelligent production in the Industrial Metaverse, then 6G and BC provide the necessary low-latency and trustworthy interaction architecture.
  • 6G: Metaverse companies use 6G networks to send and receive massive data on their Industrial Metaverse platforms. For instance, when an alert is received at the mission control center, operators can use 6G connections to remotely view and operate robots via cameras. Researchers at Nokia Bell Labs believe that compared to 5G, 6G will halve average power consumption and support a 10-fold increase in peak capacity.
  • BC: Lastly, BC, as a trustworthy distributed architecture, effectively enables trusted interactions between multiple enterprise platforms, as well as identity authentication and rights confirmation, which are indispensable in the Industrial Metaverse.
In addition, we have investigated the latest advances in theoretical research and the development of Industrial Metaverse platforms, as well as the application of each enabling technology in different phases of industrial production. In a short summary section for each technology, we briefly summarize some of the issues that still need to be addressed when integrating these technologies with the Metaverse, as well as the technical requirements in industrial scenarios.
@@
Firstly, the data input layer comprises various IoT devices used to collect data sources in the factory, such as cameras, sensors (e.g., acoustic/light/humidity sensors), and user-facing human-machine interaction devices like VR headsets and mobile phones. The collected industrial data includes text data, image data, video data, and other modalities. These multimodal data sources are transmitted to the data processing system for preprocessing, including data cleansing, integration, and transformation. They are stored in the BC through a BC gateway for pre-broadcasting. On the other hand, workers in the Industrial Metaverse can initiate data upload requests themselves. The BC packs and encapsulates the validated data through smart contracts, generating new blocks for continuous updates. When worker submits a data analysis request, they first need to send the request to the administrator. After preliminary authentication by the administrator, the request message is forwarded to the AI component for data analysis[16]. Meanwhile, the BC is responsible for recording transactions. The feedback from the data analysis is then used to update the DT. The DT subsequently sends it to the XR layer for 3D modeling, including modeling of production devices and environmental elements. The decisions obtained are displayed through XR headsets and serve various industrial scenarios in the industrial application layer, including product testing, skills training, and manufacturing. It is worth noting that the data used to update and adjust the DT model can come from the BC or directly from industrial equipment sensors, internal databases, or manually entered by operators. Through the DT component, the real world can be accurately replicated in digital form on multiple levels of virtual planes and can dynamically interact and synchronize with virtual devices on the virtual plane.
The communication between these layers and components is based on 6G, LoRa and WiFi communication network to provide the required QoS for the exchanged data. LoRa is a low-power wide area network technology specialized in long range communication for IoT devices. In the Industrial Metaverse, LoRa can support the connection of sensors and IoT devices to the virtual environment. It can be used to monitor and collect data in the real world and transmit this data to the Industrial Metaverse for users to analyze and utilize.
@@ II-A1 Research Advances of Blockchain Enabling Industrial Metaverse
The work presented in [19] introduces a communication paradigm for the native Metaverse, where users can access using encrypted addresses. A dedicated frequency band in the regional spectrum is utilized for the transmission of BC links in public domains. The controller can categorize access to specific resources to eliminate integration barriers between the Metaverse and inter-chain links. However, this encrypted address does not provide precise network topology information, and industrial IoT devices may possess mobility features, leading to a constantly changing topology structure that makes it impossible to determine the connection mode between nodes. Another practical concern is that profit-driven manufacturers typically consider whether the additional benefits of industrial synergy can cover participation costs. A potential solution is to design incentive mechanisms to attract companies to participate in routing-related activities. To address these two issues, the authors of [20] propose a layered BC architecture with a main chain and multiple sub-chains for sensor data analysis, combined with federated learning (FL) and cross-chain technology. Additionally, an age-based smart contract is utilized as an incentive mechanism to achieve trustworthy model aggregation and updates. Regarding the architecture design for dynamic topology scenarios, the authors of [21] adopt another approach for the architecture design, based on MEC. This solution allows MEC servers to possess additional resources for resource sharing and optimized utilization in non-trusted environments. This framework takes into account the heterogeneity of available server resources and user task requests, and utilizes a task assignment mechanism for edge resource sharing. This BC-based MEC deployment scheme can significantly enhance the utilization of computing resources while also ensuring system security.
@@
Numerous studies have examined distinct scenarios for amalgamating the Metaverse with BC, such as architectural modeling and remote healthcare. Though these scenarios may not be relevant to our current discussion, they may potentially be customized to the industrial production domain, such as in the modeling of production devices and remote equipment control. For instance, Building Information Modeling (BIM) technology, which is utilized in 3D architectural modeling, can be exploited through parameterization methods to facilitate speedy data computation and storage for project management processes. After exploring the components of BIM, the application of the Metaverse in virtual world construction, and the latest BC research, the authors of[22] put forward a preliminary BC-based BIM data trading framework. Moreover, the author also accentuates several potential future research topics, such as immersive design experiences in the Metaverse, the evolution, and retrospection of architectural scenarios. Bhattacharya et al.[23] raise the issue of security attacks in remote surgery, which can result in the uncontrollability of the mobile robot arm. Gupta et al.[24] initially employ the hyper-ledger fabric channel in remote surgery. In subsequent work, the author further established a BC-based collaborative access control system, which comprises three parts: the master domain, the Metaverse control, and the subordinate domain, and all metadata is encapsulated in the ledger. Thus, the operator’s behavior in the Metaverse can be recorded and utilized to enhance the robot’s behavior feedback.

### v2

@@ Abstract
As an emerging technology that enables seamless integration between the physical and virtual worlds, the Metaverse has great potential to be deployed in the industrial production field with the development of extended reality (XR) and next-generation communication networks. The Industrial Metaverse is used for product design, production operations, quality inspection, and testing. However, there is limited understanding of the enabling technologies associated with it, including the specific industrial scenarios targeted by each technology and the potential migration of technologies from other domains to the industrial sector. This paper provides a comprehensive survey of the latest literature on the Industrial Metaverse. We first analyze its advantages for industrial production, then review key enabling technologies such as blockchain (BC), privacy-preserving computing (PPC), digital twin (DT), fifth/sixth generation mobile communication technology (5G/6G), XR, and artificial intelligence (AI), and explore how these technologies support different aspects of industrial production. We also present major challenges in the Industrial Metaverse, including privacy and security concerns, resource limitations, and interoperability constraints, along with existing solutions. Finally, we outline several open issues and future research directions.
@@
Metaverse, industrial, blockchain, privacy-preserving computing, digital twin, fifth/sixth generation mobile communication technology, extended reality, artificial intelligence.
@@ I Introduction
The Metaverse is a nascent concept that presents a model of interconnection between the virtual and real worlds[1]. This digital system integrates both worlds into a cohesive system, providing users with a high degree of freedom to create content and modify modules as they wish, while also allowing different participants to define their own expressions. Regarding the conceptual description of the Metaverse, some scholars argue that “the Metaverse is a collection of virtual spaces, constructed by a series of technologies including virtual reality (VR) and augmented reality (AR). People continuously create new content through smart wearable devices as input terminals.” Others describe it as “the latest stage of visual immersion technology, essentially serving as an online digital space, applied to various fields of production and life.” Although the term “Metaverse” does not yet have a precise definition in the industry, its fundamental technological framework has gradually matured and has been successfully applied in various fields, including education, healthcare, and industrial manufacturing. The Industrial Metaverse refers to the creation of shared virtual spaces enabled by Metaverse technologies to support multi-user, multi-device industrial scenarios for 3D modeling and immersive interaction.
@@
Particularly, during the COVID-19 pandemic, many employees were forced to work from home. At the same time, continuous advancements in digital technologies have driven transformation. Many traditionally offline production activities have shifted online, allowing employees to operate remotely, similar to the concept of Digital Twin (DT) technology. DT creates virtual replicas of physical objects, systems, or processes that reflect the real-world status in real-time for monitoring, optimization, and prediction. These virtual models are tightly coupled with the real world, capable of simulating and providing feedback on the physical world. Unlike the simple mapping of DTs, the Industrial Metaverse is not merely a one-to-one replica of the real world into the virtual world. It also enables interactions and experiences that have not yet been realized in the real world. In the Industrial Metaverse, all virtual worlds and derived elements, along with digital entities in virtual spaces, can achieve online sharing, enhanced perception, and content generation. These virtual worlds, through highly immersive interactions, support the visualization, simulation, and optimization of industrial operations, allowing experiments and operations that cannot be replicated or surpassed in the real world to be conducted in a virtual environment. This provides richer and more efficient industrial application scenarios. For example, product performance can be tested through fault simulations at almost zero cost, creating experimental conditions that are impossible to replicate in reality, and extending operational capabilities through the linkage between the virtual and physical worlds. Compared to typical virtual worlds, the focus of the industrial virtual world lies in simulating industrial processes, manufacturing environments, or operational scenarios in the real world. The Industrial Metaverse places high demands on cost control in industrial manufacturing, requiring its design and implementation to consider cost efficiency and optimization, aiming to minimize deployment and maintenance costs. Additionally, the Industrial Metaverse involves collaboration between different enterprises in areas such as industrial production and supply chain management, requiring a high level of trust in the execution of interactions. Table I lists the relationships between DT, Metaverse, and Industrial Metaverse. The key benefits of Metaverse are highlighted as follows:
@@
  • Real-Time Participation through Avatars: Human participants in industrial manufacturing processes are represented by corresponding virtual avatars in the Industrial Metaverse. These avatars ensure system stability and further optimize production processes through immersive observation and sensory feedback[2].
@@
TABLE I: Comparison of DT, Metaverse, and Industrial Metaverse.
Aspect Scope Key Features Real-time Requirement User Interactivity
DT Simulation and predictive maintenance. Optimize the performance of physical assets or systems through real-time data synchronization and simulation. High. Lower.
Metaverse Entertainment, social and digital economy, etc. Focus on immersive experience and social interaction. Medium. Very high.
Industrial Metaverse Industrial application. In industrial operations, it supports data-driven decision-making and system optimization. Very high. Very high.
@@
TABLE II: Key Abbreviations.
@@
AI Artificial Intelligence
AIGC AI Generated Content
AOI Age of Information
AR Augmented Reality
ASP AI Service Provider
@@
BIM Building Information Modeling
CeAI Centralized AI
CMS Cyber-Manufacturing Systems
DeAI Decentralized AI
DP differential privacy
@@
FL Federated Learning
GANs Generative Adversarial Networks
HCI Human-computer Interaction
HE homomorphic encryption
HMDs Head-mounted Displays
IIoT Industrial Internet of Things
@@
LPWA Low-Power Wide-Area
LSTM Long Short-Term Memory
MEC Mobile Edge Computing
@@
NFT Non-Fungible Token
O&M Operations & Maintenance
PLC Product Lifecycle
PPC Privacy-Preserving Computing
@@
QoS Quality of Service
SemCom Semantic Communication
TEE Trusted Execution Environment
URLLC Ultra-reliable and Low-latency Communication
VAEs Variational Autoencoders
VR Virtual Reality
@@
XR Extended Reality
XRI XR-IoT
@@ I-A Related Works
Neal Stephenson’s novel “Snow Crash” was the first to introduce the concept of the Metaverse, a digital landscape constructed through AR technology. Since then, there has been an explosion in literature on the Metaverse, with much of it focused on its fundamental characteristics[3], empowering technologies[4], and privacy and security aspects. Some works also analyze the platform deployment and necessary components required for its application scenarios. Our research on the industrial applications of the Metaverse is comprehensive and detailed in Table III.
@@
Although some of the works do not specifically focus on industrial applications of the Metaverse, the enabling technologies involved can be directly transferred to the manufacturing industry. For instance, Dionisio et al. [3] provide a comprehensive summary of the four key characteristics of the Metaverse, namely, realism, ubiquity, interoperability, and scalability, which can be leveraged to improve manufacturing operations. Furthermore, Kye et al.[9] conduct a thorough analysis of the potential, as well as the constraints, of utilizing the Metaverse in educational contexts. This encompasses the domains of curriculum design and platform development, which could be effectively applied to skill training scenarios within the industry.
@@ I-B Comparison and Our Contributions
TABLE III: Survey papers of Industrial Metaverse.
Ref. Year Scope Technologies Key Contents
Chang et al.[11] 2022 Education, product testing, telecommuting, production optimization, smart cities Computer vision, DT, BC Three new Metaverse architectures at the edge, and technologies that help the Metaverse interact and share digital data.
Said et al.[12] 2023 None XR This work illustrates the complexity of Metaverse online learning from the perspective of HCI.
Dong et al.[14] 2023 Task-oriented communications SemCom Specific requirements for the three main types of tasks in the Industrial Metaverse.
Our paper 2023 Industry 5.0 (Industrial data collection and storage, product design, operation training, system manufacturing, industrial quality control, etc.) BC, DT, 5G/6G, XR, AI An extensive survey of the Industrial Metaverse. Particularly, For the first time, we extensively discussed the role of several key enabling technologies at different stages of industry. We comprehensively analyze the main existing problems and solutions in the Industrial Metaverse. We first emphasize more comprehensive open challenges and research directions for the characteristics of industrial scenarios.
@@ I-C Structure of The Survey
The sections of this paper are structured as follows. Section I presents an overview of the fundamental principles underlying the Metaverse and elucidates the rationale behind its integration into industrial contexts. Additionally, it provides an inventory of pertinent Metaverse literature reviews, comparing them to identify the innovative aspects of our current undertaking. Section II discusses around the five core enabling technologies of the Metaverse and summarizes the industrial applications and existing problems of each technology. Section III succinctly summarizes the four primary challenges encountered in the Industrial Metaverse. In addition, in Section IV we collate some of the existing standardization efforts and platforms for the Industrial Metaverse. Moving forward, Section V offers a comprehensive compilation of various unresolved hurdles prevailing in the contemporary Industrial Metaverse. Finally, our discourse culminates with a conclusive remark in Section VI. Table II indicates the list of abbreviations employed in this paper. Fig. 2 outlines the organization of the survey.
Figure 3: Industrial Metaverse Architecture. It consists of three layers: the data input layer, the enabling layer, and the industrial application layer. The enabling layer includes six components: AI, DT, BC, XR, the Metaverse management center, and the data processing system. In this context, the data input layer is composed of various industrial IoT devices. The collected multimodal industrial data is transmitted to the data processing system for preprocessing. The processed data can be stored in the BC or streamed to the AI component for analysis. The results of data analysis are used to respond to requests from the Metaverse management center, and decision outcomes are fed back to the BC. The BC performs updates on the network and DT. The DT achieves proportional replication and synchronization of virtual and physical devices and transfers parameter information to XR for 3D modeling. XR displays decision solutions through human-machine interaction devices and is applied in various industrial scenarios. Among them, PPC ensures that data is not disclosed during processing and analysis.
@@ II The enabling technologies for Metaverse
This section introduces several core enabling technologies for the Industrial Metaverse, including BC, PPC, DT, XR, AI, and 6G. The primary roles of these enabling technologies in the Industrial Metaverse are as follows:
@@
If DT, XR, and AI are prerequisites for enterprise decision-making and intelligent production in the Industrial Metaverse, then 5G/6G, BC, and privacy-preserving computing (PPC) provide the necessary low-latency and trustworthy interaction architecture.
  • 5G/6G: Metaverse companies use 5G/6G networks to send and receive massive data on their Industrial Metaverse platforms. For instance, when an alert is received at the mission control center, operators can use 5G/6G connections to remotely view and operate robots via cameras. Furthermore, researchers at Nokia Bell Labs believe that compared to 5G, 6G will halve average power consumption and support a 10-fold increase in peak capacity.
  • BC: As a trusted distributed architecture, BC effectively enables trusted interaction among multiple enterprise platforms and facilitates identity authentication and rights assertion when several stakeholders (such as manufacturers, suppliers, and logistics providers) need to share data without a single party having control. Compared to traditional (distributed) databases combined with centralized management, BC may offer significant advantages, and in many Industrial Metaverse scenarios, it is also indispensable.
  • PPC: Lastly, at the data level, PPC technologies such as federated learning, differential privacy (DP), and homomorphic encryption (HE) can ensure that sensitive data is processed and analyzed without being disclosed. This enables multiple parties to share industrial data and collaborate on modeling and optimization while safeguarding privacy.
In addition, we have investigated the latest advances in theoretical research and the development of Industrial Metaverse platforms, as well as the application of each enabling technology in different phases of industrial production. In a short summary section for each technology, we briefly summarize some of the issues that still need to be addressed when integrating these technologies with the Metaverse, as well as the technical requirements in industrial scenarios. Table V lists recent research works focused on enabling technologies for the Industrial Metaverse.
@@
Firstly, the data input layer comprises various IoT devices used to collect data sources in the factory, such as cameras, sensors (e.g., acoustic/light/humidity sensors), and user-facing human-machine interaction devices like VR headsets and mobile phones. The collected industrial data includes text data, image data, video data, and other modalities. These multimodal data sources are transmitted to the data processing system for preprocessing, including data cleansing, integration, and transformation. They are stored in the BC through a BC gateway for pre-broadcasting. On the other hand, workers in the Industrial Metaverse can initiate data upload requests themselves. The BC packs and encapsulates the validated data through smart contracts, generating new blocks for continuous updates. When worker submits a data analysis request, they first need to send the request to the administrator. After preliminary authentication by the administrator, the request message is forwarded to the AI component for data analysis[15]. Meanwhile, the BC is responsible for recording transactions. The feedback from the data analysis is then used to update the DT. The DT subsequently sends it to the XR layer for 3D modeling, including modeling of production devices and environmental elements. The decisions obtained are displayed through XR headsets and serve various industrial scenarios in the industrial application layer, including product testing, skills training, and manufacturing. It is worth noting that the data used to update and adjust the DT model can come from the BC or directly from industrial equipment sensors, internal databases, or manually entered by operators[16]. Through the DT component, the real world can be accurately replicated in digital form on multiple levels of virtual planes and can dynamically interact and synchronize with virtual devices on the virtual plane.
PPC ensures that during data transmission and analysis, privacy is maintained, thus allowing for secure data sharing and collaborative modeling across various participants in the Industrial Metaverse. The communication between these layers and components is based on 6G, Low-Power Wide-Area (LPWA) and WiFi communication network to provide the required QoS for the exchanged data. LPWA is a low-power wide area network technology specialized in long range communication for IoT devices. In the Industrial Metaverse, LPWA can support the connection of sensors and IoT devices to the virtual environment. It can be used to monitor and collect data in the real world and transmit this data to the Industrial Metaverse for users to analyze and utilize.
@@ II-A1 Research Advances of Blockchain Enabling Industrial Metaverse
The work presented in [19] introduces a communication paradigm for the native Metaverse, where users can access using encrypted addresses. A dedicated frequency band in the regional spectrum is utilized for the transmission of BC links in public domains. The controller can categorize access to specific resources to eliminate integration barriers between the Metaverse and inter-chain links. However, this encrypted address does not provide precise network topology information, and industrial IoT devices may possess mobility features, leading to a constantly changing topology structure that makes it impossible to determine the connection mode between nodes. Another practical concern is that profit-driven manufacturers typically consider whether the additional benefits of industrial synergy can cover participation costs. A potential solution is to design incentive mechanisms to attract companies to participate in routing-related activities. To address these two issues, the authors of [20] propose a layered BC architecture with a main chain and multiple sub-chains for sensor data analysis, combined with federated learning (FL) and cross-chain technology. Additionally, an age-based smart contract is utilized as an incentive mechanism to achieve trustworthy model aggregation and updates. Regarding the architecture design for dynamic topology scenarios, the authors of [21] adopt another approach for the architecture design, based on MEC. This solution allows MEC servers to possess additional resources for resource sharing and optimized utilization in non-trusted environments. This framework takes into account the heterogeneity of available server resources and user task requests, and utilizes a task assignment mechanism for edge resource sharing. This BC-based MEC deployment scheme can significantly enhance the utilization of computing resources while also ensuring system security[22].
@@
Numerous studies have examined distinct scenarios for amalgamating the Metaverse with BC, such as architectural modeling and remote healthcare. Though these scenarios may not be relevant to our current discussion, they may potentially be customized to the industrial production domain, such as in the modeling of production devices and remote equipment control. For instance, Building Information Modeling (BIM) technology, which is utilized in 3D architectural modeling, can be exploited through parameterization methods to facilitate speedy data computation and storage for project management processes. After exploring the components of BIM, the application of the Metaverse in virtual world construction, and the latest BC research, the authors of[23] put forward a preliminary BC-based BIM data trading framework. Moreover, the author also accentuates several potential future research topics, such as immersive design experiences in the Metaverse, the evolution, and retrospection of architectural scenarios. Bhattacharya et al.[24] raise the issue of security attacks in remote surgery, which can result in the uncontrollability of the mobile robot arm.